40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1166     Entry time: Tue Dec 2 17:56:56 2008
Author: Alberto, Rana 
Type: Configuration 
Category: PSL 
Subject: MC Alignment 
In the attempt to maximize the Mode Cleaner transmission and minimize the reflection from the steering mirrors of the MC periscope, we could not get more ~2 V at the MC Trans PD and ~ 0.5 V at MC REFL_DC. As it turned out from the SUS Drift Monitor, the reason was that the MC optics had been somehow displaced from the optimal position.

After restoring the reference position values for the mirrors and tweaking again the periscope, we got ~3V at the MC TransPD and 0.5V at the reflection.
The beam was then probably clipped at the REFL PD so that we had to adjust the alignment of one of the BS in the transmitted beam path on the AS table.
We also zeroed the WFS PDs, but not before reducing the power from the MZ, for their QPDs not to saturate.

After relocking, the transmission was 3V and the reflection ~0.3V.

The beam isnow centered on the Trans PD and REFL PD and the Mode Cleaner locked. More details on the procedure will follow.
ELOG V3.1.3-